OpenStudy (howard-wolowitz):

Can someone show me how to do this:

OpenStudy (butterfield1215):

i cant se what it says

OpenStudy (anonymous):

When you click on the image zoom in

OpenStudy (butterfield1215):

opposite over hypotenuse is how you find it

geerky42 (geerky42):

\[\sec\theta ~~=~~ \dfrac{1}{\cos\theta} ~~=~~ \dfrac{1}{\dfrac{\text{Adj}}{\text{Hyp}}}~~=~~\dfrac{\text{Hyp}}{\text{Adj}}\] So \(\sec\theta\) is ratio of Hypotenuse to Adjacent.

geerky42 (geerky42):

Do you know what value of \(c\) is?

OpenStudy (howard-wolowitz):

no i dont

OpenStudy (howard-wolowitz):

cos you mean

geerky42 (geerky42):

no, side c.

OpenStudy (howard-wolowitz):

possibly 10

geerky42 (geerky42):

we know a=b=5 Notice that is 45 45 90 triangle?

geerky42 (geerky42):

Or maybe try use Pythagorean Theorem here? \[a^2+b^2=c^2\]

OpenStudy (howard-wolowitz):

hmm ok all I have to do then is get C

geerky42 (geerky42):

kind of, you need to find \(\dfrac{c}{b}\). But we don't know \(c\). so we need to work on that.

geerky42 (geerky42):

Hyp is c,and Adj is b.

geerky42 (geerky42):

So using Pythagorean Theorem, what is c? @Howard-Wolowitz

OpenStudy (howard-wolowitz):

2 something

OpenStudy (howard-wolowitz):

I still dont get the whole C thing

OpenStudy (howard-wolowitz):

@geerky42 you cant o this without knowing C .. at least I dont think

OpenStudy (welshfella):

c^2 = 5^2 + 5^2 c^2 = 50

geerky42 (geerky42):

^